Hi! :)
I'm having a little trouble with a few questions, and FOILing. I have to convert each equation from vertex form to standard form.
1. y=(x-3)²+1
2. y=2(x-1)²-3
3. y=-3(x+4)²+1
Thanks! C:
~ Cella

Answer by stanbon(75887)

convert each equation from vertex form to standard form.
1. y=(x-3)²+1
y = x^2-6x+9+1 = x^2-6x+10
-------
2. y=2(x-1)²-3 = 2[x^2-2x+1]-3 = 2x^2-4x+2-3 = 2x^2-4x-1
-------------
3. y=-3(x+4)²+1 = -3(x^2+8x+16)+1 = -3x^2-24x-48+1 = -3x^2-24x-47
=====================
Cheers,
Stan H.
====================